# RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation

### v1

@@ Abstract
Code generation has attracted increasing attention with the rise of Large Language Models (LLMs). Many studies have developed powerful code LLMs by synthesizing code-related instruction data and applying supervised fine-tuning. However, these methods are limited by teacher model distillation and ignore the potential of iterative refinement by self-generated code. In this paper, we propose Adaptive Critique Refinement (ACR), which enables the model to refine itself by self-generated code and external critique, rather than directly imitating the code responses of the teacher model. Concretely, ACR includes a composite scoring system with LLM-as-a-Judge to evaluate the quality of code responses and a selective critique strategy with LLM-as-a-Critic to critique self-generated low-quality code responses. We develop the RefineCoder series by iteratively applying ACR, achieving continuous performance improvement on multiple code generation benchmarks. Compared to the baselines of the same size, our proposed RefineCoder series can achieve comparable or even superior performance using less data.
RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation
@@ 1 Introduction
Code generation, also called program synthesis, is a key application area of Large Language Models (LLMs) and has attracted significant attention from the research community Gulwani et al. (2017); Chen et al. (2021). Numerous studies focus on developing code-specific pre-trained models, such as CodeLlama Roziere et al. (2023), DeepSeek-Coder Guo et al. (2024), and Qwen-Coder Hui et al. (2024). These LLMs demonstrate strong capabilities in code understanding owing to their pre-training on large-scale code corpus. However, they must undergo post-training to become user-friendly and effective for code-related tasks.
@@
Figure 1: Comparison of two fine-tuning paradigms. DFT uses code distilled by teacher model, whereas ACR adaptively constructs distinct data formats infusing self-generated code and external critique.
@@
As a popular post-training technology, Distillation-based Fine-Tuning (DFT) leverages code instruction data synthesized by powerful teacher models (often proprietary LLMs like GPT-4) to fine-tune code LLMs, as shown in Figure 1. Many works Zhang et al. (2024b) based on this paradigm have been proposed recently. For example, Code Alpaca Chaudhary (2023) initially establishes this synthesis paradigm using Self-Instruct Wang et al. (2023). Subsequently, Code-related Evol-Instruct Luo et al. (2024) prompts LLMs to evolve more intricate code instructions, while OSS-Instruct Wei et al. (2024b) generates more realistic code instructions based on open-source code snippets from The Stack Kocetkov et al. (2022). In contrast, OpenCodeInterpreter Zheng et al. (2024a) enables code LLMs to learn from compiler diagnostics and human feedback by constructing multi-turn feedback data. These diverse synthesis strategies have significantly advanced the development of code LLMs. However, this paradigm of teacher model distillation inevitably suffers from a critical limitation: the performance of the student model largely relies on the teacher model, ignoring the potential of iterative refinement by self-generated code.
@@
Apart from the research on code instruction fine-tuning, some studies Chen et al. (2023); Huang et al. (2023a); Chen et al. (2024, 2025) have explored the ability of LLMs to refine self-generated code during inference. For example, Self-Debugging Chen et al. (2024) iteratively debugs self-generated code using feedback from the code executor. However, these methods that freeze parameters are essentially a form of prompt engineering with external feedback. They cannot improve the intrinsic one-off code generation ability of LLMs. Besides, multiple calls to LLMs increase the inference latency.
In this paper, we propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm to enhance the intrinsic one-shot code generation ability of LLMs, and on the basis of which we develop a series of code LLMs named RefineCoder. ACR goes beyond traditional teacher distillation by using both self-generated code and external critique to refine the model. Specifically, ACR tailors the distinct new fine-tuning dataset by first scoring and critiquing the self-generated multiple code responses, and then adaptively forms new samples according to the scoring and critiquing results. The two types of samples are shown in Figure 1. Based on the new dataset, we fine-tune RefineCoder models. Finally, we iteratively applying the above ACR process to achieve continuous improvement in code generation capabilities of RefineCoder. Notably, in ACR, we design a composite scoring system and a selective critiquing strategy to better score code responses and selectively critique low-quality code. The former combines LLM-as-a-Judge, an Elo rating mechanism Elo and Sloan (1978), and a code executor for accurate and comprehensive evaluation. The latter adaptively constructs two types of data by comparing the scores of self-generated code with those of original code. When self-generated code performs worse than the original code, an LLM-as-a-Critic is introduced to provide specific critiques.
@@
We conduct iterative ACR based on DS-Coder-6.7B-Base Guo et al. (2024) and Qwen2.5-Coder-7B-Base Hui et al. (2024), resulting in the RefineCoder series. After three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve average pass@1 improvements of 2.4 and 3.0, respectively, on the HumanEval (+), MBPP (+), LiveCodeBench, and the BigCodeBench-hard benchmark Chen et al. (2021); Austin et al. (2021); Liu et al. (2023); Jain et al. (2024); Zhuo et al. (2024). The key contributions of this paper are as follows:
@@
1) We propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm that refines code LLMs with self-generated code and external critique, on the basis of which we develop a series of strong code LLMs (RefineCoder).
@@
2) To ensure the efficacy of ACR, we design a composite scoring system with LLM-as-a-Judge and a selective critique strategy with LLM-as-a-Critic to score and critique self-generated code.
@@
3) Experimental results from the RefineCoder series show that iterative ACR continuously enhances code generation performance. After three iterations, the RefineCoder series achieves comparable or even superior performance than baselines of the same size while requiring less data.
@@ 2 Related Work
Figure 2: Overview of ACR in the $t$-th iteration. (1) Sampling: The model $M_{t}$ samples $n$ code responses $\{y^{i}\}_{i=1}^{n}$ with high temperature for a programming instruction $x$ in the dataset $D_{t}$. (2) Ranking: A composite scoring system scores the $n+1$ code responses (including $y^{0}$ from origin dataset) and ranks them, identifying a winner $y^{w}$ and $n$ losers $\{y^{l_{i}}\}_{i=1}^{n}$. (3) Refining: Depending on the identity of the $y^{w}$, the selective critique strategy constructs a new single-round data $(x,y^{w})$ or two-round critique data $(x,y^{l_{1}},c,y^{w})$, where $y^{l_{1}}$ is the highest-scoring loser, $c$ is critique. The original data $(x,y)$ is replaced with new data and all the new data form a new dataset $D_{t+1}$. (4) Training: The new model $M_{t+1}$ is fine-tuned using $D_{t+1}$.
@@
Iterative Self-Refinement of LLMs Iterative self-refinement refers to LLMs enhancing themselves iteratively by utilizing self-generated responses with the help of external signals. One line of research Huang et al. (2023b); Madaan et al. (2023); Hu et al. (2024) focuses on enabling self-correction during the inference stage by iteratively calling LLMs and incorporating external signals. In the code domain, CodeT Chen et al. (2023), Self-Debugging Chen et al. (2024), and LDB Zhong et al. (2024) follow this approach. However, these prompt engineering methods with external feedback cannot improve the intrinsic capabilities. Another line of research focuses on iteratively training the model using self-generated outputs to enhance its intrinsic capabilities Dong et al. (2025); Yuan et al. (2024); Kim et al. (2025). These methods typically rely on preference learning, such as DPO Rafailov et al. (2023). CodeLutra Tao et al. (2024) has successfully applied this approach to the code domain, but it heavily depends on golden labels, which limits its applicability. In contrast to the works above, we propose the ACR method, which achieves iterative self-refinement by only using a simple Supervised Fine-Tuning (SFT) loss. This approach is orthogonal to prompt engineering methods like Self-Debugging and is more efficient and generalizable than CodeLutra.
@@ 3.1 Overview
Adaptive Critique Refinement (ACR) aims to improve code LLMs by refining existing dataset with self-generated code and external critiques. The iterative application of ACR facilitates continuous improvement in code generation performance. As illustrated in Figure 2, ACR starts with an existing dataset and a code model fine-tuned on it. The method updates the existing dataset through a process of sampling, ranking, and refining. The updated dataset is used for the next round of fine-tuning, iteratively improving the model’s code generation capabilities.
In the following sections, we will introduce two key components of the aforementioned pipeline: the composite scoring system for ranking and the selective critiquing strategy for data refining. Finally, we will introduce the iterative training and the resulting RefineCoder model.
@@ 3.2 Composite Scoring System with LLM-as-a-Judge
Drawing on the judgment and evaluation abilities of LLMs Dong et al. (2025); Li et al. (2024), we employ an LLM-as-a-Judge as the backbone of the scoring system, supplemented by the Elo mechanism Elo and Sloan (1978) and a code executor to ensure effective evaluation. Specifically, given a programming question $x$ from the existing dataset $D_{t}$ and the corresponding code response list $\{y^{i}\}_{i=0}^{n}$, the Judge assigns pairwise relative scores for all code response pairs by identifying their differences and evaluating these discrepancies:
$\textrm{score}_{judge}^{ij},\ \textrm{score}_{judge}^{ji}=\textrm{Judge}(x,y^{i},y^{j}),$ (1)
where $i,j\in[0,n]$, $i\neq j$, and $\text{score}_{judge}^{ij}\in[1,10]$ denotes the relative score of the $i$-th response when compared to the $j$-th response. The prompt for LLM-as-a-Judge is shown in Figure 7. Next, we use Elo mechanism, a ranking algorithm that adjusts scores based on pairwise comparisons, to compute the pointwise score of each code response:
$\textrm{score}_{judge}^{i}=\textrm{Elo}\left(\left\{\textrm{score}_{judge}^{ij}\mid j\in[0,n],j\neq i\right\}\right),$ (2)
where $i\in[0,n]$ and $\textrm{score}_{judge}^{i}\in[0,1]$. For more details on Elo in the appendix A. In addition to the model scores provided by Judge and Elo, we also use a code executor to evaluate the executability of the code:
$\textrm{score}_{exec}^{i}=\textrm{Executor}(y^{i}),$ (3)
where $i\in[0,n]$ and $\text{score}_{exec}^{i}\in\{0,1\}$. Finally, we rank all code responses based on the composite score of them:
$\begin{split}&\textrm{score}^{i}=\textrm{score}^{i}_{judge}+\textrm{score}^{i}_{exec},\\ \end{split}$ (4)
$\begin{split}&y^{w},y^{l1},\dots,y^{l_{n}}=\textrm{Sort}(\{y^{i}\}_{i=0}^{n}\mid\{\textrm{score}^{i}\}_{i=0}^{n}),\end{split}$ (5)
where $\textrm{score}^{i}\in[0,2]$, $y^{w}$ and $y^{l_{1}}$ denote the code responses with the highest and second-highest scores, respectively.
@@ 3.3 Selective Critiquing Strategy with LLM-as-a-Critic
After obtaining the sorted code responses, we design the selective critiquing strategy as data construction engine of ACR. It effectively utilizes high-quality and low-quality self-generated code in distinct ways, while further providing external critiques for the low-quality data. This strategy resembles the process of a student first solving problems and then critiquing their answers based on provided solutions, rather than directly imitating the solutions Wang et al. (2025).
Concretely, when all self-generated code responses are of lower quality than the original code in the existing dataset, namely $y^{w}$ is $y^{0}$ and $y^{l_{1}}$ is the self-generated code, LLM-as-a-Critic is used to critique $y^{l_{1}}$, using $y^{0}$ as a reference (The prompt is shown in Figure 8):
$c=\textrm{Critic}(x,y^{w},y^{l_{1}}).$ (6)
Otherwise, if $y^{w}$ is the self-generated code response, we directly construct new single-round data. The data update rules are as follows:
@@
If the data to be updated comes from a two-round critique rather than a single-round one, only the instruction and the final code response will be considered. Selective refining strategy updates all data to comprise a new instruction dataset $D_{t+1}$.
@@ 3.4 Iterative Training
The previous sections describes the process of a single ACR in Figure 2. Next, we describe the iterative ACR process and define RefineCoder:
$\circ$ RefineCoder Iter0 ($M_{0}$): The code model trained using the initial SFT dataset $D_{0}$.
$\circ$ RefineCoder Iter1 ($M_{1}$): The code model trained using dataset $D_{1}$, where $D_{1}$ is obtained by refining $D_{0}$ using $M_{0}$.
$\circ$ RefineCoder Itert ($M_{t}$): The code model trained using dataset $D_{t}$, where $D_{t}$ is obtained by refining $D_{t-1}$ using $M_{t-1}$.
We fine-tune the original base model from scratch in each iteration to prevent overfitting, consistent with previous iterative training works Zelikman et al. (2022); Wang et al. (2024a); Dong et al. (2025).

### v2

# RefineCoder: Iterative Improving Large Language Models through Adaptive Critique Refinement for Code Generation

@@ Abstract
Code generation has attracted increasing attention with the rise of Large Language Models (LLMs). Many studies have developed powerful code LLMs by synthesizing code-related instruction data and applying Supervised Fine-Tuning (SFT). However, these methods are limited by teacher model distillation and ignore the potential of iterative refinement by self-generated code. In this paper, we propose Adaptive Critique Refinement (ACR), which enables the model to refine itself by self-generated code response and teacher-generated critique, rather than directly imitating the teacher-generated code responses. Concretely, ACR includes a composite scoring system with LLM-as-a-Judge to evaluate the quality of code responses and a selective critiquing strategy with LLM-as-a-Critic to critique self-generated low-quality code responses. We develop the RefineCoder series by iteratively applying ACR, achieving continuous performance improvement on multiple code generation benchmarks. Compared to the baselines of the same size, our proposed RefineCoder series can achieve superior performance using less data.
RefineCoder: Iterative Improving Large Language Models through Adaptive Critique Refinement for Code Generation
@@ 1 Introduction
Code generation, also called program synthesis, is a key application area of Large Language Models (LLMs) and has attracted significant attention from the research community Gulwani et al. (2017); Chen et al. (2021). Numerous studies focus on developing code-specific pre-trained models, such as CodeLlama Roziere et al. (2023), DeepSeek-Coder Guo et al. (2024), and Qwen-Coder Hui et al. (2024). Despite their strong capabilities in code understanding, these LLMs typically require additional post-training to become effective and user-friendly for code-related tasks.
@@
Figure 1: Comparison of two fine-tuning paradigms. DFT uses code distilled by teacher model, whereas ACR adaptively constructs distinct data formats infusing self-generated code and teacher-generated critique.
@@
As a popular post-training technology, Distillation-based Fine-Tuning (DFT) leverages code instruction data synthesized by powerful teacher models (often proprietary LLMs like GPT-4) to fine-tune code LLMs, as shown in Figure 1. Many works Zhang et al. (2024b) based on this paradigm have been proposed recently. For example, Code Alpaca Chaudhary (2023) first establishes this synthesis paradigm using Self-Instruct Wang et al. (2023). Subsequently, Code-related Evol-Instruct Luo et al. (2024) prompts LLMs to evolve more intricate code instructions, while OSS-Instruct Wei et al. (2024b) generates more realistic code instructions based on open-source code snippets Kocetkov et al. (2022). In contrast, OpenCodeInterpreter Zheng et al. (2024a) enables code LLMs to learn from compiler diagnostics and human feedback by constructing multi-turn feedback data. These diverse synthesis strategies have significantly advanced the development of code LLMs. However, this paradigm of teacher model distillation inevitably suffers from a critical limitation: the performance of the student model largely relies on the teacher model, ignoring the potential of iterative refinement by self-generated code.
@@
Apart from research on code instruction fine-tuning, some other studies Chen et al. (2023, 2024, 2025); Jiang et al. (2025) have explored the ability of LLMs to refine self-generated code during inference. For example, Self-Debugging Chen et al. (2024) iteratively debugs self-generated code using feedback from the code executor. However, these methods that freeze parameters are essentially a form of prompt engineering with external feedback. They cannot improve the intrinsic one-off code generation ability of LLMs. Besides, multiple calls to LLMs increase inference latency.
In this paper, we propose a novel fine-tuning paradigm called Adaptive Critique Refinement (ACR). It evolves through a combination of self-generated code and teacher-driven critique, thereby enhancing its capacity for one-shot code generation and surpassing the limitations of conventional teacher distillation methods. Specifically, as shown in Figure 1, ACR evaluates both self-generated and teacher-generated code, adaptively critiques low-quality self-generated responses, and generates new samples based on the evaluation and critique results. This process is analogous to a student initially attempting to solve problems independently and later comparing their solutions against given answers, rather than simply copying the provided solutions. By iteratively applying the ACR process, the model achieves continuous improvement in its code generation capabilities. Notably, we design a composite scoring system and a selective critique strategy to perform scoring and critique, respectively. The scoring system combines LLM-as-a-Judge and a code executor to provide accurate and comprehensive evaluations. The critique strategy adaptively constructs two types of data and incorporates LLM-as-a-Critic.
@@
We conduct iterative ACR based on DS-Coder-6.7B-Base Guo et al. (2024) and Qwen2.5-Coder-7B-Base Hui et al. (2024), leading to the development of the RefineCoder series. After three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B demonstrate significant improvements in average pass@1 performance on the HumanEval(+), MBPP(+), LiveCodeBench, and BigCodeBench (hard) Chen et al. (2021); Austin et al. (2021); Liu et al. (2023); Jain et al. (2024); Zhuo et al. (2024). The key contributions of this paper are as follows:
@@
1) We propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm that refines code LLMs using self-generated code and external critique, on the basis of which we develop a series of strong code LLMs namely RefineCoder.
@@
2) To ensure the effectiveness of ACR, we design a composite scoring system with LLM-as-a-Judge and a selective critiquing strategy with LLM-as-a-Critic, enabling both accurate scoring and critique of code responses.
@@
3) Experimental results from the RefineCoder series show that iterative ACR continuously improves code generation performance. After three iterations, the RefineCoder models outperform baselines of the same size while requiring less data.
@@ 2 Related Work
Figure 2: Overview of ACR in the $t$-th iteration. (1) Sampling: The model $M_{t}$ samples $n$ code responses $\{y^{i}\}_{i=1}^{n}$ with high temperature for a programming instruction $x$ in the dataset $D_{t}$. (2) Ranking: A composite scoring system first scores all self-generated code responses $\{y^{i}\}_{i=1}^{n}$, then selects the highest-scoring response for pairwise comparison with the original code response $y^{0}$ to identify the winner $y^{w}$ and the loser $y^{l_{1}}$. (3) Refining: Depending on the identity of the $y^{w}$, the selective critiquing strategy constructs a new single-turn data $(x,y^{w})$ or two-turn critique data $(x,y^{l_{1}},c,y^{w})$, where $c$ is critique. The original data $(x,y)$ is replaced with new data and all the new data form a new dataset $D_{t+1}$. (4) Training: The new model $M_{t+1}$ is fine-tuned using $D_{t+1}$.
@@
Iterative Self-Refinement of LLMs Iterative self-refinement refers to LLMs enhancing themselves iteratively by utilizing self-generated responses with the help of external signals. One line of research Huang et al. (2023); Madaan et al. (2023); Hu et al. (2024) focuses on enabling self-correction during the inference stage by iteratively calling LLMs and incorporating external signals. In the code domain, CodeT Chen et al. (2023), Self-Debugging Chen et al. (2024), and LDB Zhong et al. (2024) follow this approach. However, these prompt engineering methods with external feedback cannot improve the intrinsic capabilities. Another line of research focuses on iteratively training the model using self-generated outputs to enhance its intrinsic capabilities Dong et al. (2025); Yuan et al. (2024); Yan et al. (2025); Kim et al. (2025). These methods typically rely on preference learning, such as DPO Rafailov et al. (2023). CodeLutra Tao et al. (2024) has successfully applied this approach to the code domain, but it heavily depends on golden labels, which limits its applicability. In contrast to the works above, we propose the ACR method, which achieves iterative self-refinement by only using a simple Supervised Fine-Tuning (SFT) loss. Our approach is orthogonal to prompt engineering methods like Self-Debugging and is more efficient and generalizable than CodeLutra.
@@ 3.1 Overview
Adaptive Critique Refinement (ACR) aims to improve code LLMs by refining the existing dataset with self-generated code and external critique. The iterative application of ACR facilitates continuous improvement in code generation performance. As illustrated in Figure 2, ACR starts with a current dataset and a code model fine-tuned on it. The method updates this dataset through a process of sampling, ranking, and refining. The updated dataset is used for the next round of fine-tuning, iteratively improving the model’s code generation capabilities.
In the following sections, we will introduce two key components of the pipeline above: the composite scoring system for ranking and the selective critiquing strategy for data refining. Finally we describe the iterative training and the resulting RefineCoder model.
@@ 3.2 Composite Scoring System with LLM-as-a-Judge
Drawing on the judgment and evaluation abilities of LLMs Dong et al. (2025); Li et al. (2024), we employ LLM-as-a-Judge as the backbone of the scoring system, supplemented by a code executor to ensure effective evaluation. Specifically, we perform a multi-faceted, point-wise judgment of each self-generated code response $y^{i}$ ($1\leq i\leq n$):
$s^{i}_{1},s^{i}_{2},...,s^{i}_{m}=\textrm{Multi-Faceted Judge}(x,y^{i}),$ (1)
where $s^{i}_{j}$ denotes the score for $y^{i}$ under the $j$-th evaluation criterion, and $m$ is the total number of criteria (e.g., code correctness, clarity). Next, we calculate the final point-wise score for each self-generated code response by computing the weighted sum and normalizing the result:
$s^{i}$ $=\sum_{j=1}^{m}\alpha_{j}s_{j}^{i},$ (2)
$\text{score}^{i}_{\text{point}}$ $=\frac{s^{i}-s_{\min}}{s_{\max}-s_{\min}},$ (3)
where $\alpha_{j}$ is the weight of the $j$-th criterion, and $s_{\min}$ and $s_{\max}$ are the minimum and maximum scores across self-generated code responses. In addition, we use a code executor to evaluate the executability of all code responses (including $y^{0}$):
$\textrm{score}_{\text{exec}}^{i}=\textrm{Executor}(y^{i}),$ (4)
where $i\in[0,n]$ and $\textrm{score}_{\text{exec}}^{i}\in\{0,1\}$. The highest-scoring self-generated code is obtained by summing the point-wise score and execution score:
$y^{self}=\arg\max_{i}(\text{score}^{i}_{\text{point}}+\text{score}^{i}_{\text{exec}}).$ (5)
where $i\in[1,n]$. Finally, we determine the winner between the best self-generated code $y^{self}$ and the teacher-generated code $y^{0}$ based on the pairwise judgment and execution scores:
$\textrm{score}_{\text{pair}}^{self},\textrm{score}_{\text{pair}}^{0}=\textrm{PairJudge}(x,y^{self},y^{0}),$ (6)
$\textrm{score}^{i}=\textrm{score}^{i}_{\text{pair}}+\textrm{score}^{i}_{\text{exec}},$ (7)
$y^{w},y^{l_{1}}=\begin{cases}y^{self},y^{0},&\text{if }\textrm{score}^{self}>\textrm{score}^{0},\\ y^{0},y^{self},&\text{if }\textrm{score}^{0}>\textrm{score}^{self},\end{cases}$ (8)
where $i\in\{self,0\}$, $\textrm{score}^{i}_{pair}\in[0,1]$, and $y^{w}$ and $y^{l_{1}}$ denote the winner and loser, respectively. The prompts for point-wise and pair-wise judgment are shown in Figures 8 and 9.
@@ 3.3 Selective Critiquing Strategy with LLM-as-a-Critic
After selecting winner and loser based on quality, we design the selective critiquing strategy as the data construction engine of ACR. It effectively utilizes high-quality and low-quality self-generated code in distinct ways, while further providing external critique for the low-quality data. This strategy mirrors the process of a student first solving problems and then critiquing their answers based on provided solutions rather than directly imitating the solutions Wang et al. (2025).
Concretely, when all self-generated code responses are of lower quality than the original code in the current dataset (i.e., $y^{w}$ is $y^{0}$ from origin dataset and $y^{l_{1}}$ is the best self-generated code response), LLM-as-a-Critic is used to critique $y^{l_{1}}$, using $y^{0}$ as a reference (The prompt is shown in Figure 10):
$c=\textrm{Critic}(x,y^{w},y^{l_{1}}),$ (9)
where $c$ is the critique of $y^{l_{1}}$. Otherwise, if $y^{w}$ is the self-generated code response, we directly construct new single-turn data. The data update rules are as follows:
@@
If the data to be updated comes from a two-turn data rather than a single-turn one, only the instruction and the final code response will be considered. The selective critiquing strategy updates all data to comprise a new instruction dataset $D_{t+1}$.
@@ 3.4 Iterative Training
The previous sections describe the process of a single ACR in Figure 2. Next, we define the RefineCoder in the iterative ACR process:
$\circ$ RefineCoder Iter0 ($M_{0}$): The code model is trained using the initial SFT dataset $D_{0}$.
$\circ$ RefineCoder Iter1 ($M_{1}$): The code model is trained using dataset $D_{1}$, where $D_{1}$ is obtained by refining $D_{0}$ using $M_{0}$.
$\circ$ RefineCoder Itert ($M_{t}$): The code model is trained using dataset $D_{t}$, where $D_{t}$ is obtained by refining $D_{t-1}$ using $M_{t-1}$.
We fine-tune the original pre-trained model from scratch in each iteration to prevent overfitting, consistent with previous iterative training works Wang et al. (2024a); Dong et al. (2025).
Models Data Size LiveCodeBench BigCodeBench-hard Average
v3-v5 v5-v6 Instruct Complete
[Uncaptioned image] DSCoder-6.7B-Base / / / / 13.5 /
DSCoder-6.7B-Instruct 2BT 12.8 14.0 10.1 15.5 13.1
MagiCoder-S-DS-6.7B 185K 13.4 14.6 13.5 12.8 13.6
OpenCodeInterpreter-DS-6.7B 178K 7.3 15.2 13.5 16.9 13.2
WaveCoder-Ultra-6.7B 130K 12.9 14.5 12.8 16.9 14.3
RefineCoder-DS-6.7B   Iter0 80K 13.1 15.5 18.2 10.1 14.2
RefineCoder-DS-6.7B   Iter1 80K 14.3 17.2 18.9 14.2 16.2↑2.0
RefineCoder-DS-6.7B   Iter2 80K 14.3 17.8 19.6 14.2 16.5↑2.3
RefineCoder-DS-6.7B   Iter3 80K 14.2 17.9 20.1 15.5 16.9↑2.7
[Uncaptioned image]Qwen2.5-Coder-7B-Base / / / / 16.2    /
Qwen2.5-Coder-7B-Instruct Millions 18.1 18.9 20.3 20.3 19.4
RefineCoder-QW-7B   Iter0 80K 18.0 18.7 18.5 19.3 18.6
RefineCoder-QW-7B   Iter1 80K 19.6 21.3 20.7 21.8 20.9↑2.3
RefineCoder-QW-7B   Iter2 80K 20.5 21.5 20.6 22.9 21.4↑2.8
RefineCoder-QW-7B   Iter3 80K 21.7 21.9 20.8 22.3 21.7↑3.1
Table 1: Results on LiveCodeBench and BigCodeBench-hard. v3-v5 and v5-v6 refer to the questions from the periods 202407-202502 and 202502-202505, respectively. [Uncaptioned image] and [Uncaptioned image] denotes the two base models, and the rest are the instruction models. 2BT denotes 2B tokens.